OpenStudy (anonymous):

Consider the waves represented by the following functions. (a) y = 4 sin (3x − 15t) (b) y = 6 cos (3x + 15t − 2) (c) y = 8 sin (2x + 15t) (d) y = 8 cos (4x + 20t) (e) y = 7 sin (6x − 24t) Rank the waves represented by the above functions from the largest to the smallest according to the following conditions. (Use only ">" or "=" symbols. Do not include any spaces or parenthesis around the letters or symbols.) (i) their amplitudes (ii) their wavelengths (iii) their frequencies (iv) their periods (v) their speeds

OpenStudy (anonymous):

(1)--c=d>e>b>a (2)--c>b=a>d>e (3)--e>a>b=c>d (4)--d>c=b>a>e (5)--a>e>b=d>c :)

OpenStudy (vincent-lyon.fr):

(3) an (4) are not correct. (5) would be correct if velocity was concerned, but "speed" implies a different answer.

OpenStudy (anonymous):

yes for (5) you r correct but (3) and(4) are correct as it just depends uppon the coefficient of x:)

OpenStudy (vincent-lyon.fr):

Both depend only on the coefficient of t :)

OpenStudy (anonymous):

no as standard one is A cos(kx-wt) where k=2*pi/wavelength andwavelength=v/f=velocity/frequencyand period is inversly proportional to frequency:)

OpenStudy (vincent-lyon.fr):

This logical chain is wrong because when you write λ=v/f, you are assuming that all waves travel with the same velocity. It is simpler: ω = 2π/T , so period and frequency depend only on ω.

OpenStudy (anonymous):

oh i just forgot to take that into consideration my fault sry:)

OpenStudy (vincent-lyon.fr):

yw :))